# Dynamic Pre-training: Towards Efficient and Scalable All-in-One Image Restoration

### v1

@@ 6 Conclusion
Supplemental Material
Here, we have discussed the details about the transformer block used in our DyNet, additional sample images from our Million-IRD dataset, additional visual results comparison between PromptIR [36] and the proposed DyNet-L in all-in-one setting.
Figure S1: Overview of the Transformer block used in our DyNet network. The Transformer block is composed of two sub-modules, the Multi Dconv head transposed attention module (MDTA) and the Gated Dconv feed-forward network (GDFN).
A Transformer Block Details
@@
B Visual Results Comparison
We show additional visual results comparison between PromptIR [36] and our DyNet-L under from all-in-one setting.
Figure S2: Comparative analysis of image denoising by all-in-one methods on the BSD68 dataset [30]. Our DyNet-L reduces noise, producing more sharp and clear image compared to the PromptIR [36].
Figure S3: Comparative analysis of image deraining by all-in-one methods on the Rain100L dataset [16]. Our DyNet-L successfully eliminates rain streaks, producing clear, rain-free images as compared to the recent PromptIR [36].
Figure S4: Comparative analysis of image dehazing by all-in-one methods on the SOTS dataset [25]. Our approach reduces haze, producing more clear image compared to the PromptIR [36].
C Breakdown of images in Our Million-IRD Dataset
We combine the existing high-quality, high-resolution natural image datasets such as LSDIR [27], DIV2K [2], Flickr2K [35], and NTIRE [4]. Collectively these datasets have 90K images having spatial size ranging between $<1024^{2},4096^{2}>$. Their breakdown is given in Table S1. We also show the additional sample images from our Million-IRD dataset in Fig. S5 and Fig. S6.
Table S1: Overview of the number of images curated from each database for our Million-IRD dataset.
Dataset NTIRE [4] DIV2K [2] Flikr2K [35] LSDIR [27] Laion-HR [42]
@@ C Breakdown of images in Our Million-IRD Dataset
Figure S5: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs. This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc.
Figure S6: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs. This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc.

### v2

@@ 6 Conclusion
Appendix A Appendix (Supplementary Material)
Here, we have discussed the details about the transformer block used in our DyNet, additional sample images from our Million-IRD dataset, additional visual results comparison between PromptIR (Potlapalli et al., 2023) and the proposed DyNet in all-in-one setting.
Figure 8: Overview of the Transformer block used in our DyNet network. The Transformer block is composed of two sub-modules, the Multi Dconv head transposed attention module (MDTA) and the Gated Dconv feed-forward network (GDFN).
Appendix B Transformer Block Details
@@
Appendix C Visual Results Comparison
We show additional visual results comparison between PromptIR (Potlapalli et al., 2023) and our DyNet under all-in-one setting.
Figure 9: Comparative analysis of image denoising by all-in-one methods on the BSD68 dataset (Martin et al., 2001) and Urban100 (Huang et al., 2015). Our DyNet reduces noise, producing more sharper and clearer image compared to the PromptIR (Potlapalli et al., 2023).
Figure 10: Comparative analysis of image deraining by all-in-one methods on the Rain100L dataset (Fan et al., 2019). Our DyNet-L successfully eliminates rain streaks, producing clear, rain-free images as compared to the recent PromptIR Potlapalli et al. (2023).
Figure 11: Comparative analysis of image dehazing by all-in-one methods on the SOTS dataset (Li et al., 2018). Our approach reduces haze, producing more clear image compared to the PromptIR (Potlapalli et al., 2023).
Appendix D Breakdown of images in Our Million-IRD Dataset
We combine the existing high-quality, high-resolution natural image datasets such as LSDIR (Li et al., 2023a), DIV2K (Agustsson & Timofte, 2017), Flickr2K (Online, ), and NTIRE (Ancuti et al., 2021). Collectively these datasets have 90K images having spatial size ranging between $<1024^{2},4096^{2}>$. Their breakdown is given in Table 7. We also show the additional sample images from our Million-IRD dataset in Fig. 12 and Fig. 13.
Table 7: Overview of the number of images curated from each database for our Million-IRD dataset.
Dataset NTIRE DIV2K Flikr2K LSDIR Laion-HR
 (Ancuti et al., 2021)  (Agustsson & Timofte, 2017)  (Online, )  (Li et al., 2023a)  (Schuhmann et al., 2022)
@@ Appendix D Breakdown of images in Our Million-IRD Dataset
Figure 12: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs. This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc.
Figure 13: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs. This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc.